wiki:UC20190805

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 8/05/19

Attending: Mark Klaene (APO), Bill Ketzeback (APO), Kevin Schlaufman (JHU), Ben Williams (UW), Kal Kadlec (UW), John Wisniewski (OU), Anne Verbiscer (UVa), Eric Bellm (UW), Joanne Hughes Clark (Seattle U), John Bally (CU), Sarah Tuttle (UW), Nancy Chanover (NMSU), Misty Bentz (GSU)


User feedback and comments from institutional representatives:

  • Wyoming - no report
  • Washington - nothing to report
  • UVa - nothing to report
  • Seattle - nothing to report
  • Oklahoma -nothing to report
  • NMSU - nothing to report
  • NAPG -no report
  • JHU - nothing to report
  • Georgia State - nothing to report
  • FGCU - no report
  • Colorado - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 6/27/2019 - 8/01/2019

0) Overview

The main highlight of activity has been around the summer shutdown. As noted it was unusually short this year, in part because we lost our aluminizing slot at KPNO, and also because we had an extra-long shutdown last year due to M2 and M3 work. The new M2 and M3 overcoats reduces the need for us to recoat them over the summer, and a lot of instrument issues that transpired over the year kept us from having to do major instrument servicing this summer. A lot of IT work was done, primarily due to the change in IT staff over the past year. We also installed new fiber runs from the computer room to the telescope.

1) Telescope

The telescope has been performing nominally after a small glitch coming out of shutdown during engineering, which affected the primary mirror support structure (PMSS) and motor preloads. M1, M2 and M3 were washed and the telescope was recollimated. Several other PM items were accomplished.

2) Instruments

DIS: The DIS blue camera continues to see contamination issues reappear rapidly after servicing while the red camera performance is deteriorating more slowly. Thus there are no plans to service DIS blue at this time. DIS red got a little worse but is still good. This was not serviced over summer shutdown.

TripleSpec is operational.

Agile is operational.

Echelle is operational. The new Ecam is in place and operable. The IOL is steadily getting worse; we are monitoring it closely to determine when and if we will intervene.

NIC-FPS is operational.

ARCTIC is operational; we are working on some minor filter issues.

3) ARCSAT

Operating nominally. The new guide scope mount has too much flexure so a new one is being designed.


Additional telescope and instrument discussion:

The main activities on the mountain over the past month were summer shutdown and preparation for the KOSMOS commissioning. This was a very short summer shutdown, which was intentional (partly because of the extra-long one last year, and also because we didn’t have recoat smaller mirrors and we didn’t get a time slot in which to recoat the primary mirror at KPNO). All mirrors got washed. M1 still has cosmetic issues from bats/birds/moths, but these do not impact the overall throughput losses very much. The windowing capability is now working in E-cam. The ARCES interorder light is beginning to level out a bit; it had gotten worse throughout the spring but we may see an improvement as the temperatures cool off moving into fall. If not we may consider a pumping or at least a warming servicing of it. Not much was done with instrumentation because they had been serviced throughout the year. Previous experience has shown that when we do service the instruments in the summer during warm and humid conditions we find that we have to do it again shortly thereafter, so we're trying to break that cycle this time. With the short summer shutdown we have had a few nights or partial nights of summer observing, which is unusual. We still have several weeks of monsoons left to go.

Bill has a new design for the ARCSAT guider scope that is ~85% complete. Bill and Adam Kowalski (CU) worked on an alternative approach for getting high cadence images with FlareCam that involves bypassing APC and using a script that runs under MaxIm. It works well for someone observing the same target for many hours but it requires more intervention from the obs-spec to use it so it is not at the point of being broadly available.


Q3 3.5m and ARCSAT scheduling:

There is still some 3.5m time available in Q3: DD01 on Aug 22B, 29B (dark), and Sep 05B, and OPEN Aug 09B (requested) and 15A. [UPDATE: both Aug 09B and 15A are now spoken for. The late August/early September DD01 time remains available.] For ARCSAT, there are still several open weeks in late August and early/mid September. Please submit requests for any open Q3 time to Nancy, Ben, Russet, and cc your institutional scheduler.


Q4 3.5m and ARCSAT scheduling:

All allocations were sent out and we are expecting to receive back all of the cover sheets by Aug 27. We modified the proposal form to make it more clear that users can ask for item in units of hours instead of half nights. Please make sure your department members are using the newest version of the form, which is available here https://www.apo.nmsu.edu/35m_operations/35m_manual/General_Information/proposal_form.txt Please also indicate if any early Q4 requests could be executed in September in the event that the KOSMOS commissioning window slips into October. Once the Q4 schedule for the 3.5m is complete (early September) we will send out a call for Q4 ARCSAT proposals.


NA2 Baffle Project:

Work is proceeding. The photonics firm doing the stray light analysis determined that the telescope is generally well baffled. Mechanical drawings for fabrication of the new secondary baffle are about to get started.


KOSMOS Update:

The UW TEG now has most of the parts in house and they are moving into the assembly and testing phase. Kal successfully removed the detector from the KOSMOS dewar for installation into the COSMOS dewar; one cable needs to be rewired and then they can continue testing. They've done some basic testing within the lab of the slit viewer. There are quite a few slits of different sizes. Conor is making a big push on the software so by the time they get the CCD running they hope to be able to talk to it via the new software. They are still finalizing the schedule for the pre-ship review and getting the instrument to the observatory.

  • Question: what spectral resolutions are available? Answer: there are only 2 grisms - one red, one blue - with R ~ 2200 for an 0.8” slit (so resolution will scale as go to larger slit).
  • Question: what is the process for getting few additional grisms? Answer: we should survey the user community to determine what is needed/desired to help prioritize such purchases. Nancy and Misty will work on this.
  • Question: how hard is it to change out grisms? Answer: There is a grism wheel with 6 positions, we would generally leave one clear so in principle we could mount 3-4 other “standard” grisms and then have other one-offs that could be installed/removed as needed. They do not need to be red-blue pairs; we can install whatever we want. The wheel is easily accessible and this should not be too onerous.
  • Question: what is the clear slot used for? Answer: for field acquisition (but we wouldn’t technically need it since we will have a new guide camera).
  • Question: what is the FOV in acquisition mode? Answer: 5' x 4.3’.

Response to Call for White Papers for 2020 Decadal Survey

The white paper was submitted; thanks to those who submitted comments. It is one of ~ 330 APC white papers (+ ~640 science white papers!). There were 4 others that were closely related to ours, which Nancy summarized:

  1. Stephen Ridgway (NOAO): Ground Based Optical Astronomy — Keeping the Innovation Window Open (advocates for open access to private facilities)
  2. Sally Oey (University of Michigan): 2020 Vision: Towards a Sustainable OIR System (advocates for continued access to and investment in modest aperture telescopes)
  3. Joan Najita (NOAO): Investing for Discovery in Astronomy (there are diverse paths to discovery, small teams disrupt/large teams develop, results from NOAO community survey in May 2019: Respondents were asked to indicate the importance of each resource in a funding-limited environment. The highest priority is (1) observing time, followed by (2) archival data products, (3) new instrumentation, (4) software and data pipelines, (5) student or mentor training, and (6) new telescopes.)
  4. Lori Allen (NOAO): The NOAO Mid-Scale Observatories (current state of affairs at KPNO and assets in Chile, possible upgrades to modest telescopes, AEON promotion

To make it more broadly accessible Nancy will post it on the wiki (unless there are a lot of APC white papers on astro-ph). Nancy will discuss with Mike Crenshaw whether the ARC Board should be notified of the existence of this white paper prior to posting it.


Discussion of Near-Term Action Items from Strategic Planning Meeting

  • Getting DAWG underway: Nancy is finalizing organizational structure and charge to group
  • Modifications to proposal form - we want to make it easier and streamline the process for triggering ToOs; now people should contact Ben (daytime) or Russet (evening/night) if they want to trigger. We are still working on streamlining this process further.
  • Nancy requested that everyone send her citations for any publications from the past year. The FINAL DEADLINE is Wednesday August 7, 2019.
  • Nancy is exploring the idea of hiring an undergraduate from NMSU's Journalism and Mass Communications department as a science communications intern.
  • It was decided not to invite the media to APO on the night of July 20 (the Apollo 11 50th anniversary) because although APOLLO observations were scheduled it was uncertain whether we would be back on sky that night after summer shutdown. Instead, Nancy tried to work with NMSU Communications on a press release about it but there were challenges getting all of the necessary information in time. Instead, there were several local events held (public talks by Katie, Karen, and Russet), so the contributions of APO were advertised that way.
  • Upcoming:
    • The Jan. 2020 AAS abstract deadline is in mid-October; Nancy would like input regarding what sort of presence we should have at the AAS meeting (booth? poster? special session?)
    • Several class trips are scheduled for this fall, so Nancy would like to revisit the idea of having the students do a video hackathon (students filming themselves) to create some training videos. Nancy will talk to several people to try and identify a few starter topics that they can try out.

Data Analysis Working Group (DAWG)

Nancy is working to identify a chair and to finalize the initial charge to the group.


ACTION ITEMS:

Open action items from previous meetings:

  • Ben will send out a link to KOSMOS info for interested prospective users. STATUS: OPEN. This is intended to be a page where prospective users could learn about the general capabilities of KOSMOS at a glance. Sarah and Kal have been worrying more about getting KOSMOS ready for commissioning, so this remains open. Sarah envisions this to be an entry point with capabilities that will be built up to a full manual that they can link to the APO web site. Ben will work with Kal to give him examples of what we're looking for. STATUS: OPEN. There is a draft available and several people are reviewing it to assess whether it contains all of the information that a prospective KOSMOS user would be looking for.
  • All UC reps should notify their users that if their Q3 proposals require DIS Blue for their science, they should contact Russet ASAP and let her know. STATUS: CLOSED.
  • UC reps should let users know that the problem with headers has now been fixed, but it affected all data between June 7-July 1. Nancy will send a note to 35m-general about this. STATUS: CLOSED.
  • UC reps should mention as their users to be on the lookout for echelle interorder light issues, particularly if doing emission line science. STATUS: CLOSED.
  • UC reps should notify users of any Q3 OPEN/DD time remaining. STATUS: OPEN.
  • UC reps should encourage people to take a look at the Astro2020 white paper and contribute. STATUS: CLOSED.
  • Nancy will post the minutes from the Strategic Planning meeting to the users' wiki. STATUS: CLOSED.
  • Nancy will contact current DAWG members to finalize the organization of this working group. STATUS: CLOSED.

New action items from this meeting:

  • Nancy will coordinate with Mike Crenshaw regarding distribution of the Astro2020 APC White Paper.
  • Nancy and Misty will survey the 3.5m user community regarding any potential additional grisms that would be desired for KOSMOS.
  • Nancy will identify some possible topics for training video hacks to be done by visiting classes of students.
  • Anyone with thoughts on what sort of presence we should have at AAS should send them to Nancy.

All Other Business

None.


Next meeting:

The next Users Committee telecon will be on Tuesday September 3 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.